OpenStudy (anonymous):

factor 324= x^3-12x^2+27x-324 PLEASE HELP

OpenStudy (anonymous):

0 = x^3 - 12x^2 + 27x - 324 flip the second and third terms 0 = x^3 + 27x - 12x^2 - 324 group the first two terms and the last two terms 0 = (x^3 + 27x) - (12x^2 + 324) factor out common terms 0 = x(x^2 + 27) - 12(x^2 + 27) apply the Distributive Property 0 = (x - 12)(x^2 + 27) We have two factors. Set each equal to zero and solve for x. x - 12 = 0 x = +12 x^2 + 27 = 0 x^2 = -27 x = √(-27) x = √(-1) * √(27) x = i * √(9 * 3) x = i * (+-3√3) x = -3i√3 and +3i√3 So the three solutions are +12 and -3i√3 and +3i√3 Im not sure if this is the right answer, but you can check it.. LOL

OpenStudy (anonymous):

@Trexy I think you missed the 324 at the beginning of the question.

OpenStudy (anonymous):

ooh well.. Sorry about that..lol Can you check it @Calcmathlete ??

OpenStudy (anonymous):

lol. I did it in my head and factoring by grouping doesn't work so well...you'd have to go through rational root theorem, descartes rule of signs, etc and it's well...meh...

OpenStudy (anonymous):

i did all of that i just need to factor it

OpenStudy (anonymous):

Thank You for Correcting it @Calcmathlete XD

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Can any one please help Find the solutions to this equation algebraically using the Fundamental Theorem of Algebra, the Rational Root Theorem, Descartes' Rule of Signs, and the Factor Theorem. I did everything except the factor part what do I need to do

OpenStudy (anonymous):

It does not factor if there is a 324 on the LHS

OpenStudy (anonymous):

Here are the roots that Mathematica gave \[ \begin{array}{c} x=4+\frac{1}{3} \sqrt[3]{9018-243 \sqrt{1373}}+\sqrt[3]{334+9 \sqrt{1373}} \\ x=4-\frac{1}{6} \left(1+i \sqrt{3}\right) \sqrt[3]{9018-243 \sqrt{1373}}-\frac{1}{2} \left(1-i \sqrt{3}\right) \sqrt[3]{334+9 \sqrt{1373}} \\ x=4-\frac{1}{6} \left(1-i \sqrt{3}\right) \sqrt[3]{9018-243 \sqrt{1373}}-\frac{1}{2} \left(1+i \sqrt{3}\right) \sqrt[3]{334+9 \sqrt{1373}} \\ \end{array} \]

OpenStudy (anonymous):

It does not have rational roots. Does it really have a 324 on the LHS?

OpenStudy (anonymous):

Did you type your equation right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you can't factor it?

OpenStudy (anonymous):

Not with rational numbers.

OpenStudy (anonymous):

I think the 324 on the LHS must be zero. In the case, one of the posts above solved it for you.